What annual rate of return is earned on a 5,100 investment made in year 4 when it grows to $11,300 by the end of year 10?
Answer #1

Annual rate of return of the investment is 14.18 %

Explanation:

Total return = (Ending value – Beginning value)/Beginning value

                      = ($ 11,300 - $ 5,100)/ $ 5,100

                      = $ 6,200 /$ 5,100 = 1.2156862745098

Annual return = (1+ Total return) (1/No of years) - 1

                          = (1 + 1.2156862745098) (1/6) - 1

                          = (2.2156862745098) (1/6) - 1

                          = (2.2156862745098) 0.166666667 - 1

                          = 1.14178599406128 - 1

                          = 0.14178599406128 or 14.18 %